# Characteristics and prevalence of fake social media profiles with AI-generated faces

### v1

@@ 4.3 Heuristics for identification
During the annotation, we identified some heuristic indicators that can help spot GAN-generated profiles. The most distinctive feature lies in the flaws. Although GANs can generate realistic human faces, other parts of the images are often featured with defects. For instance, GANs sometimes have difficulty producing realistic hats (Fig. 4(a)), glasses (Fig. 4(b)), and ears (Fig. 4(d)). Similarly, users can pay attention to accessories, such as earrings (Fig. 4(d)) and necklaces, that GANs often struggle with. Clothing details, especially collars, are also problematic for GANs (Fig. 4(c)). Last but not least, the background sometimes can hint at the artificial nature of a picture (Fig. 4(b)). While annotating TwitterGAN accounts, we specifically examined the defects in their profile pictures. Two annotators agreed that 60.1% of them have clear defects (Cohen’s Kappa = 0.90).
@@ 5.1 Eye locations
Our detection method leverages the observation that GANs place the eyes in the same locations in all generated pictures, as demonstrated in Fig. 1(a,b). This feature has been reported and exploited to distinguish true and fake human faces (Yang et al., 2019b). Note that our task differs from previous methodology papers since our goal is to identify GAN-generated profiles from random Twitter profiles. As illustrated in Fig. 1(c), many profiles do not feature human faces. And for those containing authentic human faces, we hypothesize that very few of them happen to have their eyes placed in the same locations as the GAN-generated ones.
To test our hypothesis, we need to obtain the eye coordinates of input images. We test different face recognition software and find that the Python package face_recognition111111github.com/ageitgey/face_recognition provides an effective and efficient solution. The package returns an empty list if no faces are detected in the input images. For images containing faces, the package returns the coordinates of the landmarks, such as the eyes and nose, of each face (Kazemi and Sullivan, 2014). We focus on the eyes and obtain the center coordinate of each eye by averaging its landmark points. Since the profile images have different resolutions, we normalize the coordinates using the image height and width so that all the re-scaled coordinates are in the unit interval.
@@
We apply the package to the profiles in TwitterGAN, AcademicGAN, and RandomTwitter. Figure 5 shows the kernel density estimation distributions of the detected eye positions. For the GAN-generated profiles in both TwitterGAN and AcademicGAN, the eye locations are highly concentrated and identical. On the other hand, the eye locations of random profiles in RandomTwitter are scattered across the space. These findings confirm our hypothesis and indicate this feature can be leveraged to distinguish GAN-generated and random Twitter profiles.
Inspecting the results of face_recognition yields a few more critical observations. The package can identify exactly one face from all AcademicGAN images and all TwitterGAN images except for one with clear defects. On the other hand, the package can only identify faces from around 35.7% profiles in RandomTwitter. The remaining profiles either do not contain faces or are misclassified by the package. Among the profiles with detected faces, about 92.3% have one face detected, while the others have multiple faces detected. Although face_recognition is not perfect in face detection, it works very well on the GAN-generated faces, ensuring a high recall for our objective.
@@ 5.2 GANEyeDistance metric
Based on the promising results in Fig. 5, we propose a metric called GANEyeDistance to quantify the distances between the detected eyes and the expected coordinates of the eyes in GAN-generated pictures.
@@
To apply Eq. (1), we also need the ground truth values of $\vec{L}_{\text{GAN}}$ and $\vec{R}_{\text{GAN}}$. Inspecting the profiles in TwitterGAN suggests that the profile pictures are subject to modifications such as cropping and stretching, introducing noise. On the other hand, pictures in AcademicGAN have better consistency and quality, so we use the average values of their eye coordinates as our ground truth. Our implementation of GANEyeDistance is publicly available.121212github.com/osome-iu/fake_gan_accounts
@@
To demonstrate the effectiveness of GANEyeDistance, we calculate this value for every image in all three datasets and show their distributions in Fig. 6(a). We can see that the profiles in AcademicGAN and TwitterGAN have GANEyeDistance values very close to 0, as expected. The results of RandomTwitter show a bimodal distribution: the GANEyeDistance values of the profiles with one detected face are distributed across a wide range, and the remaining images have a GANEyeDistance value of 1. The difference in the distributions suggests that GANEyeDistance can effectively distinguish GAN-generated profiles from random Twitter profiles. In Fig. 6(b), we zoom in to better inspect the difference among the three datasets. The GANEyeDistance values for AcademicGAN are closest to 0. The results for TwitterGAN deviate from 0 slightly due to noise. The AcademicGAN distribution has a very small overlap with those of the GAN-generated profiles.
Despite the promising results shown in Fig. 6, we emphasize again that GANEyeDistance only indicates the distance between the locations of detected eyes and the ground truth. It cannot truly distinguish GAN-generated faces from authentic human faces. This means that although GAN-generated faces have GANEyeDistance close to zero, images with very small GANEyeDistance are not necessarily GAN-generated. In other words, a relatively loose threshold for GANEyeDistance can yield a very high recall for GAN-generated profiles but also introduce false positives.
For instance, if we use 0.02 as the threshold and consider all images with $\mathcal{G}<0.02$ as positive cases, then the recall for AcademicGAN and TwitterGAN is 100% and 99.5%, respectively. At the same time, 0.46% (1,181) of the random profiles are labeled as positive, so further inspection is needed to determine their nature.
@@ 5.3 Prevalence estimation
Two authors manually classify these profiles into three categories: (1) highly likely GAN-generated, (2) likely GAN-generated, and (3) not GAN-generated. Profiles in the first category demonstrate clear and unique features of GAN-generated human faces, such as the defects illustrated in Fig. 4. Profiles in the second category are very similar to GAN-generated faces; however, no definitive evidence can be identified.
The Cohen’s Kappa for the ratings is 0.85, indicating high agreement between the two annotators. 54 profiles are coded as highly likely generated by GANs by both annotators, and 113 are believed to be likely or highly likely generated by GANs by both annotators. Considering that these profiles are selected from the 254,275 accounts in RandomTwitter and our method might miss some positive cases, we estimate a lower bound on the prevalence of active Twitter users using GAN-generated profiles to be between 0.021–0.044%. These accounts were responsible for 0.022–0.053% of the tweets in RandomTwitter.
@@ 6 Discussion
Using a random sample, we further estimate the prevalence of such accounts on Twitter. We first build a machine learning algorithm exploiting the common feature of GAN-generated faces to narrow down the potential GAN-generated profiles, then perform manual annotation to identify those likely to be synthetic. Our results suggest that the percentage of active Twitter accounts using GAN-generated profiles is between 0.021% and 0.044%. Since the recall of our machine learning model is not 100%, and GAN-generated images that are heavily stretched or tilted can evade our method, we might miss some positive cases in our analysis. Also, the labeling criteria used in our annotation process are very conservative — we only consider a profile positive when two annotators are both highly certain about the label. Therefore, these estimates should be considered as a lower bound.
@@
To put the percentages into perspective, let us consider the total number of daily active Twitter users. Pfeffer et al. (2023) collected all 375 million tweets posted on September 21, 2022, and found 40,199,195 active users. Using this number, we estimate that at least 8,537–17,864 daily active accounts on Twitter use GAN-generated profiles.
@@
Our analysis and estimation have some limitations. We only include a few groups of accounts in TwitterGAN, so their behavior patterns might not generalize to other cases. Although we have proposed an effective method to capture more accounts using GAN-generated profiles in the wild, further analysis is not feasible since Twitter ended free access to its API. Also, since we only focus on one platform, the findings might not generalize to other platforms, such as Facebook and LinkedIn, that have been reported to suffer from the same issue. Unfortunately, limited access to their data makes it impossible for researchers to study inauthentic accounts on those platforms. The Digital Services Act by the European Union mandates that large social media platforms, including Twitter and Meta, provide data access to researchers, bringing hope.131313ec.europa.eu/commission/presscorner/detail/en/IP_23_2413 Nonetheless, it remains uncertain whether researchers can obtain the necessary data to expand on this research.
@@
The heuristics and detection methods introduced in our study do not necessarily work on other generative AI models. For example, diffusion models (Song and Ermon, 2019; Ho et al., 2020) yield better image-generation outcomes than GANs (Dhariwal and Nichol, 2021; Yang et al., 2023). This architecture has been adopted by many AI companies, including OpenAI (Ramesh et al., 2022), Midjourney,141414docs.midjourney.com and Stability AI,151515stability.ai/stable-image to create advanced multimodal models. The images generated by these models are astonishing and free from the issues of GAN-generated ones. Preliminary studies on detecting diffusion model-generated images suggest that the techniques that work on GAN-generated images are no longer effective (Ricker et al., 2022), requiring new detection paradigms.
@@
Despite these limitations, our findings provide empirical evidence of the threats posed by generative AI. We show that GAN-generated images have already been deployed at scale to create large amounts of convincing fake accounts. Similarly, researchers recently discovered a botnet that uses ChatGPT to generate human-like text and conduct realistic interactions on Twitter (Yang and Menczer, 2023). Combining these AI models can lead to more convincing and realistic fake accounts on social media. In fact, recent research suggests that advanced AI models enable the creation of autonomous agents that can ingest multimodal inputs, process the data, reason, and then leverage various tools to achieve their objectives (Xi et al., 2023). Despite the usefulness of these intelligent agents, the risk of adversarial actors abusing them is high.
@@ Acknowledgements
We thank Twitter user @conspirator0 for flagging candidate accounts using GAN-generated profile images.

### v2

@@ 4.3 Heuristics for identification
During the annotation, we identified some heuristic indicators that can help spot GAN-generated profiles. The most distinctive feature lies in the flaws. Although GANs can generate realistic human faces, other parts of the images are often featured with defects. For instance, GANs sometimes have difficulty producing realistic hats (Figure 4(a)), glasses (Figure 4(b)), and ears (Figure 4(d)). Similarly, users can pay attention to accessories, such as earrings (Figure 4(d)) and necklaces, that GANs often struggle with. Clothing details, especially collars, are also problematic for GANs (Figure 4(c)). Last but not least, the background sometimes can hint at the artificial nature of a picture (Figure 4(b)). While annotating TwitterGAN accounts, we specifically examined the defects in their profile pictures. Two annotators agreed that 60.1% of them have clear defects (Cohen’s Kappa = 0.90).
@@ 5.1 Eye locations
Our detection method leverages the observation that GANs place the eyes in the same locations in all generated pictures, as demonstrated in Figure 1(a,b). This feature has been reported and exploited to distinguish true and fake human faces (Yang et al., 2019b). Note that our task differs from previous methodology papers since our goal is to identify GAN-generated profiles from random Twitter profiles. As illustrated in Figure 1(c), many profiles do not feature human faces. And for those containing authentic human faces, we hypothesize that very few of them happen to have their eyes placed in the same locations as the GAN-generated ones.
To test our hypothesis, we need to obtain the eye coordinates of input images. We test different face recognition software and find that the Python package face_recognition [^5] https://github.com/ageitgey/face_recognition provides an effective and efficient solution. The package returns an empty list if no faces are detected in the input images. For images containing faces, the package returns the coordinates of the landmarks, such as the eyes and nose, of each face (Kazemi and Sullivan, 2014). We focus on the eyes and obtain the center coordinate of each eye by averaging its landmark points. Since the profile images have different resolutions, we normalize the coordinates using the image height and width so that all the re-scaled coordinates are in the unit interval.
@@
We apply the package to the profiles in TwitterGAN, AcademicGAN, and RandomTwitter. Figure 5 shows the kernel density estimation distributions of the detected eye positions. For the GAN-generated profiles in both TwitterGAN and AcademicGAN, the eye locations are highly concentrated and identical. On the other hand, the eye locations of random profiles in RandomTwitter are scattered across the space. These findings confirm our hypothesis and suggest that TwitterGAN images were likely generated by the same algorithm as those in AcademicGAN. In turn, this indicates eye locations can be leveraged to distinguish GAN-generated and random Twitter profiles.
Inspecting the results of face_recognition yields a few more critical observations. The package can identify exactly one face from each AcademicGAN image and each TwitterGAN image, except for one with clear defects. On the other hand, the package can only identify faces from around 35.7% profiles in RandomTwitter. The remaining profiles either do not contain faces or are misclassified by the package. Among the profiles with detected faces, about 92.3% have one face detected, while the others have multiple faces detected. Although face_recognition is not perfect in face detection, it works very well on the GAN-generated faces, ensuring a high recall for our objective.
@@ 5.2 GANEyeDistance metric
Based on the promising results in Figure 5, we propose a metric called GANEyeDistance to quantify the distances between the detected eyes and the expected coordinates of the eyes in GAN-generated pictures.
@@
To apply Eq. (1), we also need the ground truth values of $\vec{L}_{\text{GAN}}$ and $\vec{R}_{\text{GAN}}$. Inspecting the profiles in TwitterGAN suggests that the profile pictures are subject to modifications such as cropping and stretching, introducing noise. On the other hand, pictures in AcademicGAN have better consistency and quality, so we use the average values of their eye coordinates as our ground truth. Our implementation of GANEyeDistance is publicly available. [^6] https://github.com/osome-iu/fake_gan_accounts
@@
To demonstrate the effectiveness of GANEyeDistance, we calculate this value for every image in all three datasets and show their distributions in Figure 6(a). We can see that the profiles in AcademicGAN and TwitterGAN have GANEyeDistance values very close to 0, as expected. The results of RandomTwitter show a bimodal distribution: the GANEyeDistance values of the profiles with one detected face are distributed across a wide range, and the remaining images have a GANEyeDistance value of 1. The difference in the distributions suggests that GANEyeDistance can effectively distinguish GAN-generated profiles from random Twitter profiles. In Figure 6(b), we zoom in to better inspect the difference among the three datasets. The GANEyeDistance values for AcademicGAN are closest to 0. The results for TwitterGAN deviate from 0 slightly due to noise. The AcademicGAN distribution has a very small overlap with those of the GAN-generated profiles.
Despite the promising results shown in Figure 6, we emphasize again that GANEyeDistance only indicates the distance between the locations of detected eyes and the ground truth. It cannot truly distinguish GAN-generated faces from authentic human faces. This means that although GAN-generated faces have GANEyeDistance close to zero, images with very small GANEyeDistance are not necessarily GAN-generated. In other words, a relatively loose threshold for GANEyeDistance can yield a very high recall for GAN-generated profiles but also introduce false positives.
For instance, if we use 0.02 as the threshold and consider all images with $\mathcal{G}<0.02$ as positive cases, then the recall for AcademicGAN and TwitterGAN is 100% and 99.4%, respectively. At the same time, 0.46% (1,181) of the random profiles are labeled as positive, so further inspection is needed to determine their nature.
@@ 5.3 Prevalence estimation
To reduce the false positives where the algorithm flags authentic human photos as GAN-generated, two authors manually classify these profiles into three categories: (1) highly likely GAN-generated, (2) likely GAN-generated, and (3) not GAN-generated. Profiles in the first category demonstrate clear and unique features of GAN-generated human faces, such as the defects illustrated in Figure 4. Profiles in the second category are very similar to GAN-generated faces; however, no definitive evidence can be identified.
The Cohen’s Kappa for the ratings is 0.85, indicating high agreement between the two annotators. 54 profiles are coded as highly likely generated by GANs by both annotators, and 113 are believed to be likely or highly likely generated by GANs by both annotators. Profiles are considered not GAN-generated if at least one annotator believes so. Considering that these profiles are selected from the 254,275 accounts in RandomTwitter and our method might miss some positive cases, we estimate a lower bound on the prevalence of active Twitter users using GAN-generated profiles to be between 0.021–0.044%. These accounts were responsible for 0.022–0.053% of the tweets in RandomTwitter.
@@ 6 Discussion
Using a random sample, we further estimate the prevalence of such accounts on Twitter. We first build a heuristic method exploiting the common feature of GAN-generated faces to narrow down the potential GAN-generated profiles, then perform manual annotation to identify those likely to be synthetic. Our results suggest that the percentage of active Twitter accounts using GAN-generated profiles is between 0.021% and 0.044%. To put the percentages into perspective, let us consider the total number of daily active Twitter users. Pfeffer et al. (2023) collected all 375 million tweets posted on September 21, 2022, and found 40,199,195 active users. Using this number, we estimate that at least 8,537–17,864 daily active accounts on Twitter use GAN-generated profiles.
Since the recall of our machine learning model is not 100%, and GAN-generated images that are heavily stretched or tilted can evade our method, we might miss some positive cases in our analysis. Also, the labeling criteria used in our annotation process are very conservative—we only consider a profile positive when two annotators are both highly certain about the label. Therefore, these estimates should be considered as a lower bound. In fact, a recent report by Ricker et al. (2024) produced a similar but slightly higher estimate of 0.052% using different detection methods and excluding accounts with default profile images in the denominator.
@@
Our analysis and estimation have some limitations. We only include a few groups of accounts in TwitterGAN, so their behavior patterns might not generalize to other cases. Although we have proposed an effective method to capture more accounts using GAN-generated profiles in the wild, further analysis is not feasible since Twitter ended free access to its API. Also, since we only focus on one platform, the findings might not generalize to other platforms, such as Facebook and LinkedIn, that have been reported to suffer from the same issue. Unfortunately, limited access to their data makes it impossible for researchers to study inauthentic accounts on those platforms. The Digital Services Act by the European Union mandates that large social media platforms, including Twitter and Meta, provide data access to researchers, bringing hope. [^7] https://ec.europa.eu/commission/presscorner/detail/en/IP_23_2413 Nonetheless, it remains uncertain whether researchers can obtain the necessary data to expand on this research.
@@
The heuristics and detection methods introduced in our study do not necessarily work on other generative AI models. For example, diffusion models (Song and Ermon, 2019; Ho et al., 2020) yield better image-generation outcomes than GANs (Dhariwal and Nichol, 2021; Yang et al., 2023). This architecture has been adopted by many AI companies, including OpenAI (Ramesh et al., 2022), Midjourney, [^8] https://docs.midjourney.com and Stability AI, [^9] https://stability.ai/stable-image to create advanced multimodal models. The images generated by these models are astonishing and free from the issues of GAN-generated ones. Preliminary studies on detecting diffusion model-generated images suggest that the techniques that work on GAN-generated images are no longer effective (Ricker et al., 2022), requiring new detection paradigms.
@@
Despite these limitations, our findings provide empirical evidence of the threats posed by generative AI. We show that GAN-generated images have already been deployed at scale to create large amounts of convincing fake accounts. Many more social media accounts likely have already been using more advanced AI models to generate more realistic fake profiles. We even uncover a group of accounts that also leverage ChatGPT to generate human-like text and conduct realistic interactions (Yang and Menczer, 2024). Combining these AI models can lead to more convincing and realistic fake accounts on social media. In fact, recent research suggests that advanced AI models enable the creation of autonomous agents that can ingest multimodal inputs, process the data, reason, and then leverage various tools to achieve their objectives (Xi et al., 2023). Despite the usefulness of these intelligent agents, the risk of adversarial actors abusing them is high.
@@ Acknowledgements
We thank https://generated.photos for sharing their datasets for our research and Twitter user @conspirator0 for flagging candidate accounts using GAN-generated profile images.